Suppose that market demand is given by the equation qd=121.00−p, and market supply is given by the equation qs=p−16.00. If the government imposes a price ceiling on this good at a price of $30.00, what would be the change in consumer's surplus relative to the market equilibrium? When making your calculation, assume that the consumers who value the good the most are the ones who purchase the good. Also, assume that these consumers purchase the good at the ceiling price. Round your answer to two decimal places.

   


1
Expert's answer
2021-10-04T14:05:15-0400

Qd = 121 - P


Qs= P - 16


At equilibrium Qd = Qs


121 - P = P - 16


2P = 137


P = 137/2

= 68.5


Q = 121 - 68.5

= 52.5


Producer surplus = 1/2 × 68.5 × 52.5

= 1798.125


Suppose government imposes a price ceiling of $30


Qd = 121 - 30

= 91


Consumer surplus = 1/2 × 91 × 68.5

= 3116.75


Qs = 30 - 16


Producer surplus = 1/2 × 14 × 30

= 210